OpenStudy (anonymous):

How can I find the vertex in the standard form of quadratic function: y = 2x^2 + 5 ? i don't understand because there's no b value

OpenStudy (aravindg):

b=0 here

OpenStudy (kirbykirby):

WHat do you use as "b" in the standard form? What did your tell you what the standard form was with your notation?

OpenStudy (anonymous):

y-5=2x^2 if we put y-5=Y and x-0 =X we have the standard form as X^2=1/2 Y X^2=4*(1/8) Y hence reqd vertex is 0,5

OpenStudy (anonymous):

it's using the standard form f(x) = ax^2 + bx + c

OpenStudy (kirbykirby):

Ok so as @AravindG said, b=0 because ax^2+0x+c = ax^2 +c, and here a = 2, and c=5

OpenStudy (aravindg):

yep here it will be \[y=2x^2+0x+5\]

OpenStudy (anonymous):

thank you so so much

OpenStudy (aravindg):

yw :)